# A single-loop SPIDER-type stochastic subgradient method for expectation-constrained nonconvex nonsmooth optimization

### v1

@@ 3.2 Convergence analysis and complexity results
$\textstyle 2\rho\alpha_{k}\mathbb{E}\left[\left\langle\widehat{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k)},\beta{\mathbf{w}}^{(k)}\right\rangle\right]\leq\mathbb{E}\left[{\alpha_{k}\rho^{2}}\|\widehat{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k)}\|^{2}\right]+\mathbb{E}\left[{\alpha_{k}\beta^{2}}{}\|{\mathbf{w}}^{(k)}\|^{2}\right]$
$=$ $\textstyle\frac{\alpha_{k}}{4}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({{\mathbf{x}}}^{(k)})\right\|^{2}\right]+{\alpha_{k}\beta^{2}}{}\mathbb{E}\left[\left\|\left(\mathrm{Proj}_{[0,1]}\left(\frac{u^{(k)}}{\nu}\right)-\mathrm{Proj}_{[0,1]}\left(\frac{g({\mathbf{x}}^{(k)})}{\nu}\right)\right)\boldsymbol{\zeta}_{g}^{(k)}\right\|^{2}\right]$
$\leq$ $\textstyle\frac{\alpha_{k}}{4}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({{\mathbf{x}}}^{(k)})\right\|^{2}\right]+\alpha_{k}\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})\mathbb{E}\left[\left|\mathrm{Proj}_{[0,1]}\left(\frac{u^{(k)}}{\nu}\right)-\mathrm{Proj}_{[0,1]}\left(\frac{g({\mathbf{x}}^{(k)})}{\nu}\right)\right|^{2}\right]$
$\leq$ $\textstyle\frac{\alpha_{k}}{4}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({{\mathbf{x}}}^{(k)})\right\|^{2}\right]+\alpha_{k}\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})\mathbb{E}\left[\left|\frac{u^{(k)}}{\nu}-\frac{g({\mathbf{x}}^{(k)})}{\nu}\right|^{2}\right]$
$=$ $\frac{\alpha_{k}}{4}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({{\mathbf{x}}}^{(k)})\right\|^{2}\right]+\frac{\alpha_{k}\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}r_{g}^{(k)},$
where the first equality follows by Eq. 14 and Eq. 26, the second inequality comes from 1(c) and 4(b), and the third inequality uses 1-Lipschitz of Proj operator. Substituting the above inequality back into Eq. 35 yields Eq. 32.
@@
Under 1 and 4, given $\overline{\epsilon}>0$, suppose that $\{{\mathbf{x}}^{(k)}\}_{k=0}^{Kq-1}$ and $\{u^{(k)}\}_{k=0}^{Kq-1}$ are generated by Algorithm 1 with
(36) $S_{1}=\lceil 16\overline{\epsilon}^{-2}\nu^{-2}(l_{g}^{2}+\sigma_{2}^{2})\beta^{2}\sigma^{2}\rceil,\,q=S_{2}=\lceil\sqrt{S_{1}}\rceil,\,$
(37) $\alpha_{k}=\alpha=\min\left\{\frac{\overline{\epsilon}^{2}}{16\rho L_{F}^{2}},\frac{\overline{\epsilon}\nu}{4\beta\sqrt{l_{g}^{2}+\sigma_{2}^{2}}L_{g}L_{F}}\right\},\text{ and }K=\lceil 16\overline{\epsilon}^{-2}q^{-1}\alpha^{-1}\Delta\rceil.$
Here, $\Delta:=\mathbb{E}[F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(0)})]-\min_{{\mathbf{x}}}\mathbb{E}[F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}})]$ and $L_{F}^{2}=2(\sigma_{1}^{2}+l_{f}^{2}+\beta^{2}\sigma_{2}^{2}+\beta^{2}l_{g}^{2})$. It holds
@@
Since $\alpha_{k}=\alpha$Lemma 3.1 implies
$\sum_{k=0}^{Kq-1}\frac{\alpha_{k}\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}r_{g}^{(k)}\leq\frac{Kq\alpha\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\left(\frac{\alpha^{2}qL_{F}^{2}L_{g}^{2}}{S_{2}}+\frac{\sigma^{2}}{S_{1}}\right).$
Summing Eq. 32 from $k=0$ to $Kq-1$ and applying the above bound, we derive
@@
(39) $+\frac{Kq\alpha\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\left(\frac{\alpha^{2}qL_{F}^{2}L_{g}^{2}}{S_{2}}+\frac{\sigma^{2}}{S_{1}}\right)+Kq{\alpha^{2}{\rho}L_{F}^{2}}{}.$
@@
$\quad\quad\quad+\frac{4\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\left(\frac{\alpha^{2}qL_{F}^{2}L_{g}^{2}}{S_{2}}+\frac{\sigma^{2}}{S_{1}}\right)+4{\alpha{\rho}L_{F}^{2}}{}$
@@
Now we bound the RHS of Eq. 40. From $\alpha\leq\frac{\overline{\epsilon}^{2}}{16\rho L_{F}^{2}}$, it follows that $4{\alpha{\rho}L_{F}^{2}}\leq\frac{\overline{\epsilon}^{2}}{4}$. The choice $S_{1}=\lceil 16\overline{\epsilon}^{-2}\beta^{2}\nu^{-2}(l_{g}^{2}+\sigma_{2}^{2})\sigma^{2}\rceil$ yields $\frac{4\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\frac{\sigma^{2}}{S_{1}}\leq\frac{\overline{\epsilon}^{2}}{4}$. The conditions $q=S_{2}=\lceil\sqrt{S_{1}}\rceil$ and $\alpha\leq\frac{\overline{\epsilon}\nu}{4\beta\sqrt{l_{g}^{2}+\sigma_{2}^{2}}L_{g}L_{F}}$ imply $\frac{4\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\frac{\alpha^{2}qL_{F}^{2}L_{g}^{2}}{S_{2}}\leq\frac{\overline{\epsilon}^{2}}{4}$. Also, we obtain $\frac{4\Delta}{Kq\alpha}\leq\frac{\overline{\epsilon}^{2}}{4}$ from $K=\lceil 16\overline{\epsilon}^{-2}q^{-1}\alpha^{-1}\Delta\rceil$. Thus we have the desired result.
@@
Under the conditions of Corollary 3.3, given $\overline{\epsilon}>0$, let $\{{\mathbf{x}}^{(k)}\}_{k=0}^{Kq-1}$ and $\{u^{(k)}\}_{k=0}^{Kq-1}$ be generated by Algorithm 1 with $S_{1}=N$, (i.e., taking the whole data), and other parameters set to those in Eq. 36Eq. 37. Then Eq. 38 holds.
From Theorem 3.6 and Theorem 3.8, we can directly derive the complexity of Algorithm 1 to produce a near-$\overline{\epsilon}$ stationary solution of Eq. PP in expectation by choosing an iterate from $\{{\mathbf{x}}^{(k)}\}_{k=0}^{Kq-1}$ uniformly at random.
@@
Under 1 and 4Algorithm 1 can produce a near-$\overline{\epsilon}$ stationary solution of Eq. PP in expectation with $Kq=O\left(\overline{\epsilon}^{-4}\max\{1,\frac{\overline{\epsilon}}{\nu}\}\right)$ stochastic subgradients and $K(S_{1}+(q-1)S_{2})=O\left(\overline{\epsilon}^{-5}\nu^{-1}\max\{1,\frac{\overline{\epsilon}}{\nu}\}\right)$ stochastic function evaluations on $g$. If $\mathbb{P}_{g}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$, the number of stochastic function evaluations becomes $O\left(N+\sqrt{N}\overline{\epsilon}^{-4}\max\{1,\frac{\overline{\epsilon}}{\nu}\}\right)$.
3.3 Identifying an $(\epsilon,\epsilon)$-KKT point of problem Eq. P in expectation
In this subsection, we demonstrate that Algorithm 1 can find an $(\epsilon,\epsilon)$-KKT point for problem Eq. P under certain specific scenarios. The following proposition addresses the case with a convex constraint or a special weakly-convex case.
@@
[…]
@@
    $O(\epsilon^{-4})$ OGC/CGC and $O(\epsilon^{-6})$ CFC if $\mathbb{P}_{g}$ is a general distribution;
@@
    $O(\epsilon^{-4})$ OGC/CGC and $O(N+\sqrt{N}\epsilon^{-4})$ CFC if $\mathbb{P}_{g}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$.
@@
Choose $\overline{\epsilon}$, $\beta$, and $\nu$ so that the conditions in Eq. 17 hold and $\beta>~{}\frac{B_{f}+\rho_{f}D^{2}}{B_{g}-\rho_{g}D^{2}}$. In addition, set $S_{1}$ as in Eq. 36 if $\mathbb{P}_{g}$ is a general distribution and $S_{1}=N$ , (i.e., taking the whole data), if $\mathbb{P}_{g}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$. Then set $S_{2},$ $q$, $\alpha_{k}$ and $K$ to those in Eq. 36 and Eq. 37. The claims now directly follow from Theorem 2.8(i)-(ii) and Corollary 3.9.

### v2

@@ 3.2 Convergence analysis
$2\rho\alpha_{k}\mathbb{E}\left[\left\langle\widehat{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k)},\beta{\mathbf{w}}^{(k)}\right\rangle\right]\leq\mathbb{E}\left[{\alpha_{k}\rho^{2}}\|\widehat{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k)}\|^{2}\right]+\mathbb{E}\left[{\alpha_{k}\beta^{2}}{}\|{\mathbf{w}}^{(k)}\|^{2}\right]$
$=$ $\frac{\alpha_{k}}{4}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({{\mathbf{x}}}^{(k)})\right\|^{2}\right]$
$+{\alpha_{k}\beta^{2}}{}\mathbb{E}\left[\left\|\sum_{i=1}^{m}\left(\mathrm{Proj}_{[0,1]}\left(\frac{u_{i}^{(k)}}{\nu}\right)-\mathrm{Proj}_{[0,1]}\left(\frac{g_{i}({\mathbf{x}}^{(k)})}{\nu}\right)\right)\boldsymbol{\zeta}_{g_{i}}^{(k)}\right\|^{2}\right]$
$\leq$ $\frac{\alpha_{k}}{4}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({{\mathbf{x}}}^{(k)})\right\|^{2}\right]$
$+\alpha_{k}\beta^{2}m(l_{g}^{2}+\sigma_{2}^{2})\mathbb{E}\left[\sum_{i=1}^{m}\left|\mathrm{Proj}_{[0,1]}\left(\frac{u_{i}^{(k)}}{\nu}\right)-\mathrm{Proj}_{[0,1]}\left(\frac{g_{i}({\mathbf{x}}^{(k)})}{\nu}\right)\right|^{2}\right]$
$\leq$ $\frac{\alpha_{k}}{4}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({{\mathbf{x}}}^{(k)})\right\|^{2}\right]+\alpha_{k}\beta^{2}m(l_{g}^{2}+\sigma_{2}^{2})\mathbb{E}\left[\sum_{i=1}^{m}\left|\frac{u_{i}^{(k)}}{\nu}-\frac{g_{i}({\mathbf{x}}^{(k)})}{\nu}\right|^{2}\right]$
$=$ $\frac{\alpha_{k}}{4}\mathbb{E}\left[\left\|\nabla F^{\nu}_{\frac{1}{2\rho}}({{\mathbf{x}}}^{(k)})\right\|^{2}\right]+\frac{\alpha_{k}\beta^{2}m(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}r_{g}^{(k)},$
where the first equality follows from Eq. 14 and Eq. 27, and the second inequality comes from 1(c) and 3(b). Substituting the above inequality back into Eq. 36 yields Eq. 33.
@@
Under 1 and 3, given $\overline{\epsilon}>0$, suppose that $\{{\mathbf{x}}^{(k)}\}_{k=0}^{Kq-1}$ and $\{{\mathbf{u}}^{(k)}\}_{k=0}^{Kq-1}$ are generated by Algorithm 1 with
(37) $S_{1}=\lceil 16\overline{\epsilon}^{-2}\nu^{-2}m(l_{g}^{2}+\sigma_{2}^{2})\beta^{2}\sigma^{2}\rceil,\,q=S_{2}=\lceil\sqrt{S_{1}}\rceil,\,$
(38) $\alpha_{k}=\alpha=\min\Big{\{}\frac{\overline{\epsilon}^{2}}{16\rho L_{F}^{2}},\frac{\overline{\epsilon}\nu}{4\beta\sqrt{m(l_{g}^{2}+\sigma_{2}^{2})}L_{g}L_{F}}\Big{\}},\text{ and }K=\lceil 16\overline{\epsilon}^{-2}q^{-1}\alpha^{-1}\Delta\rceil.$
Here, $\Delta:=F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{(0)})-\min_{{\mathbf{x}}}F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}})$, and $L_{F}$ is given in Lemma 3.1. It holds
@@
[…]
$\sum_{k=0}^{Kq-1}\frac{\alpha_{k}\beta^{2}(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}r_{g}^{(k)}\leq\frac{Kq\alpha\beta^{2}m(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\left(\frac{\alpha^{2}qL_{F}^{2}L_{g}^{2}}{S_{2}}+\frac{\sigma^{2}}{S_{1}}\right).$
Summing Eq. 33 from $k=0$ to Kq1Kq-1 and applying the above bound, we derive
@@
(40) $+\frac{Kq\alpha\beta^{2}m(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\left(\frac{\alpha^{2}qL_{F}^{2}L_{g}^{2}}{S_{2}}+\frac{\sigma^{2}}{S_{1}}\right)+Kq{\alpha^{2}{\rho}L_{F}^{2}}{}.$
@@
$\quad\quad\quad+\frac{4\beta^{2}m(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\left(\frac{\alpha^{2}qL_{F}^{2}L_{g}^{2}}{S_{2}}+\frac{\sigma^{2}}{S_{1}}\right)+4{\alpha{\rho}L_{F}^{2}}{}$
@@
Now we bound the RHS of Eq. 41. From $\alpha\leq\frac{\overline{\epsilon}^{2}}{16\rho L_{F}^{2}}$, it follows that $4{\alpha{\rho}L_{F}^{2}}\leq\frac{\overline{\epsilon}^{2}}{4}$. The choice $S_{1}=\lceil 16\overline{\epsilon}^{-2}\beta^{2}m\nu^{-2}(l_{g}^{2}+\sigma_{2}^{2})\sigma^{2}\rceil$ yields $\frac{4\beta^{2}m(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\frac{\sigma^{2}}{S_{1}}\leq\frac{\overline{\epsilon}^{2}}{4}$. The conditions $q=S_{2}$ and $\alpha\leq\frac{\overline{\epsilon}\nu}{4\beta\sqrt{m(l_{g}^{2}+\sigma_{2}^{2})}L_{g}L_{F}}$ imply $\frac{4\beta^{2}m(l_{g}^{2}+\sigma_{2}^{2})}{\nu^{2}}\frac{\alpha^{2}qL_{F}^{2}L_{g}^{2}}{S_{2}}\leq\frac{\overline{\epsilon}^{2}}{4}$. Also, we obtain $\frac{4\Delta}{Kq\alpha}\leq\frac{\overline{\epsilon}^{2}}{4}$ from $K=\lceil 16\overline{\epsilon}^{-2}q^{-1}\alpha^{-1}\Delta\rceil$. Thus we have the desired result.
@@
Under the conditions of Corollary 3.3, given $\overline{\epsilon}>0$, let $\{{\mathbf{x}}^{(k)}\}_{k=0}^{Kq-1}$ and $\{{\mathbf{u}}^{(k)}\}_{k=0}^{Kq-1}$ be generated by Algorithm 1 with $S_{1}=N$ (i.e., taking the whole data), and other parameters set to those in Eq. 37Eq. 38. Then Eq. 39 holds.
3.3 Complexity result
In this subsection, we demonstrate that Algorithm 1 can find an $(\epsilon,\epsilon)$-KKT point for problem Eq. P, using the results in Theorem 3.6 and Theorem 2.9. From these two theorems, we can directly derive the complexity of Algorithm 1 to produce a near-$\overline{\epsilon}$ stationary solution of Eq. PP in expectation by choosing an iterate from $\{{\mathbf{x}}^{(k)}\}_{k=0}^{Kq-1}$ uniformly at random.
@@ 3.3 Complexity result
Under 1 and 3Algorithm 1 can produce a near-$\overline{\epsilon}$ stationary solution of Eq. PP in expectation with $Kq=O\left(\overline{\epsilon}^{-4}\max\{1,\frac{\overline{\epsilon}}{\nu}\}\right)$ stochastic subgradients and $K(S_{1}+(q-1)S_{2})=O\left(\overline{\epsilon}^{-5}\nu^{-1}\max\{1,\frac{\overline{\epsilon}}{\nu}\}\right)$ stochastic function evaluations on $\mathbf{g}$}. If $\mathbb{P}_{{\mathbf{g}}}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$, the number of stochastic function evaluations becomes $O\left(N+\sqrt{N}\overline{\epsilon}^{-4}\max\{1,\frac{\overline{\epsilon}}{\nu}\}\right)$.
We then present the complexity results in the following proposition.
@@
Suppose that 1, 2 with $B_{g}>m\rho_{g}D^{2}$, and 3 hold. Then, given $\epsilon>0$, Algorithm 1 can produce an $(\epsilon,\epsilon)$-KKT point of problem Eq. P in expectation with
@@
    $O(\epsilon^{-4})$ OGC/CGC and $O(\epsilon^{-6})$ CFC if $\mathbb{P}_{{\mathbf{g}}}$ is a general distribution;
@@
    $O(\epsilon^{-4})$ OGC/CGC and $O(N+\sqrt{N}\epsilon^{-4})$ CFC if $\mathbb{P}_{{\mathbf{g}}}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$.
@@
Choose $\overline{\epsilon}$, $\beta$, and $\nu$ so that the conditions in Eq. 17 hold. In addition, set $S_{1}$ as in Eq. 37 if $\mathbb{P}_{{\mathbf{g}}}$ is a general distribution and $S_{1}=N$ (i.e., taking the whole data), if $\mathbb{P}_{{\mathbf{g}}}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$. Then set $S_{2},$ $q$, $\alpha_{k}$ and $K$ to those in Eq. 37 and Eq. 38. The claims now directly follow from Theorem 2.9 and Corollary 3.9.